# A Zero-Shot Open-Vocabulary Pipeline for Dialogue Understanding

### v1

@@ 3.2 Domain Classification
As domain classification is the first step in the pipeline, its accuracy is critical for downstream tasks like slot tracking. Any misclassification can have a cascading impact. To mitigate this, we ensure that domain classification is independent for each dialogue turn. For instance, if a turn is classified as taxi-related, we don’t carry over the “taxi” to subsequent turns. To further improve the accuracy, we enforce strict guidelines in the prompts to ensure the classification remains grounded in the specific dialogue context and is not biased toward domains that might be inferred from specific phrases (the instruction part of the prompt). For example, given the utterance “I want a taxi to go to the hotel.”, the domain should be strictly classified as “taxi” rather than “hotel” and “taxi”. This is also to solve the classification of the closing turns (e.g., “have a nice day”, “anything else for today”), which contain generic terms and hurt the performance as also showed in Hudeček and Dusek (2023).
@@ 5.2 Dialogue State Tracking
Table 3: Comparison of LDST and SRP on the MultiWOZ 2.2 Test Split.
@@
Since the language model version might have a large impact on results, we compare LDST Feng et al. (2023) and SRP using the same model which is shown to be capable: GPT-3.5-Turbo. By using the same dataset, slot descriptions, possible values, and language model across different prompts (ours versus LDST), we show that SRP outperforms LDST by a large margin (9%).
@@ 6.1 Error Analysis
Figure 2: JGA Per Domain for MutliWOZ 2.4
@@ 6.2 Analysis of LLM Prompt Requests
Finally, we evaluate the efficiency and scalability of our models by analyzing the number of LLM prompt requests required, a key metric for computational cost and deployment viability. Table 4 shows the average API calls per dialogue (see App. C for calculation details and the prompt length consideration) for MultiWOZ 2.4. The results show that our QA and SRP approaches need 96.35% and 97.08% fewer requests, respectively, compared to the “All Slots” approaches, and 86.33% and 89.06% fewer requests, respectively, compared to the “Turn Domains Slots” approaches. However, compared to a baseline method that includes all schema slots for each turn in a single prompt, our approaches necessitate 120.16% and 77.02% more requests, respectively. This increase is due to a significant reduction in prompt length, as we now use only the slots of the active domains for each turn rather than listing all slots for all domains. Additionally, there is a substantial improvement in JGA when compared to the results and prompts reported by Heck et al. (2023). disable,size=,inlinedisable,size=,inlinetodo: disable,size=,inlinenot very clear to me with this much reduction in efficiency, we get this much performance improvement maybe mention small models as well disable,size=,inline, color=yellowdisable,size=,inline, color=yellowtodo: disable,size=,inline, color=yellowDone. I rephrased the paragraph. Originally. it was: However, compared to a baseline method where all schema slots for each turn are included in a single prompt, our approaches require 120.16% and 77.02% more requests, respectively. This increase can be attributed to significant impacts on prompt length and model performance when compared to the results and prompts reported by Heck et al. (2023).
@@
    Track all slots from all domains
@@
    Track slots from the turn domains only but each in a single LLM request
@@
Table 4: Comparison of Different Approaches and Their Number of LLM Prompt Requests for the MultiWOZ Test Split
@@ 7 Conclusion
In this work, we introduce a zero-shot, open-vocabulary Dialogue State Tracking (DST) system that integrates domain classification and DST in a single pipeline. By reformulating DST as a question-answering task and employing adaptable prompting techniques, our system dynamically adapts to new slot values without additional fine-tuning. We find that by selecting appropriate techniques—either QA-based or well-structured prompts tailored to the size of the language model—we can surpass SOTA models even without relying on any predefined values from ontologies. Although integrating the domain classification stage reduces the overall pipeline performance, it is essential for creating a practical system. Finally, we demonstrate the computational efficiency of our techniques by smartly selecting the slots to query and optimizing the prompts to track all input slots in one go, rather than querying the system for each slot individually.
@@ A.1 Domain Classification
For the dialogue turns between the user and system, Which service is the user asking for? Choose from the following: restaurant, attraction, hotel, taxi, train, bus, hospital, police. If you can’t find a service, return None.
@@ Appendix B Language Models Hyperparameters
Appendix C Number of LLM Prompt Requests Calculations
@@
While the prompt length is also a key metric to consider, it can be tricky to identify as it depends on several factors including the task description, the slot values, the slot description, the frequency of sending the task (for each dialogue, slot, domain), and the LLM API itself. Some LLMs require sending the conversation history each time you add a new message, whereas others do not.

### v3

@@ 3.2 Domain Classification
As domain classification is the first step in the pipeline, its accuracy is critical for downstream tasks like slot tracking. Any misclassification can have a cascading impact. To mitigate this, we ensure that domain classification is independent for each dialogue turn. For instance, if a turn is classified as taxi-related, we don’t carry over the “taxi” to subsequent turns. To further improve the accuracy, we enforce strict guidelines in the prompts (in the initial task description provided for the LLM to generate the SRP) to ensure the classification remains grounded in the specific dialogue context and is not biased toward domains that might be inferred from specific phrases. For example, given the utterance “I want a taxi to go to the hotel.”, the domain should be strictly classified as “taxi” rather than “hotel” and “taxi”. This is also to solve the classification of the closing turns (e.g., “have a nice day”, “anything else for today”), which contain generic terms and hurt the performance as also showed in Hudeček and Dusek (2023).
@@ 5.2 Dialogue State Tracking
Table 3: Comparison of LDST and SRP on the MultiWOZ 2.2 Test Split in the same test setup.
@@
Since the language model version might have a large impact on results, we compare LDST Feng et al. (2023) and SRP using the same model which is shown to be capable: GPT-3.5-Turbo. By using the same dataset, slot descriptions, possible values, and language model across different prompts (ours versus LDST), we show that SRP outperforms LDST by a large margin (9%) as shown in Table 3.
@@ 6.1 Error Analysis
Figure 2: JGA Comparison Across Domains in MultiWOZ 2.4 Dataset for GPT-4-Turbo and Llama 3 Models
@@ 6.2 Analysis of LLM Prompt Requests
Finally, we evaluate the efficiency and scalability of our models by analyzing the number of LLM prompt requests required, a key metric for computational cost and deployment viability. Table 4 shows the average API calls per dialogue (see App. C for details on LLM request metrics and calculation methods) for MultiWOZ 2.4. The results show that our QA and SRP approaches need 96.35% and 97.08% fewer requests, respectively, compared to the “All Slots” approaches, and 86.33% and 89.06% fewer requests, respectively, compared to the “Turn Domains Slots” approaches. However, compared to a baseline method that includes all schema slots for each turn in a single prompt, our approaches necessitate 120.16% and 77.02% more requests, respectively. This increase is due to a significant reduction in prompt length, as we now use only the slots of the active domains rather than all domains. Additionally, there is a substantial improvement in JGA when compared to the results and prompts reported by Heck et al. (2023). disable,size=,inlinedisable,size=,inlinetodo: disable,size=,inlinenot very clear to me with this much reduction in efficiency, we get this much performance improvement maybe mention small models as well disable,size=,inline, color=yellowdisable,size=,inline, color=yellowtodo: disable,size=,inline, color=yellowDone. I rephrased the paragraph. Originally. it was: However, compared to a baseline method where all schema slots for each turn are included in a single prompt, our approaches require 120.16% and 77.02% more requests, respectively. This increase can be attributed to significant impacts on prompt length and model performance when compared to the results and prompts reported by Heck et al. (2023).
@@
    Track all slots from all domains each in a single LLM request
@@
    Track slots from active domains and each in a single LLM request
@@
Table 4: Comparison of Different Approaches and Their Number of LLM Prompt Requests for the MultiWOZ
@@ 7 Conclusion
In this work, we introduce a zero-shot, open-vocabulary Dialogue State Tracking (DST) system that integrates domain classification and DST in a single pipeline. By reformulating DST as a question-answering task and employing adaptable prompting techniques, our system adapts to new slot values without additional fine-tuning. We find that by selecting appropriate techniques—either QA-based or well-structured prompts tailored to the size of the language model—we can surpass SOTA models even without relying on any predefined values from ontologies. Although integrating the domain classification stage reduces the pipeline performance, it is essential for creating a practical system. Finally, we demonstrate the computational efficiency of our techniques by smartly selecting the slots to query and optimizing the prompts to track all slots in one go, rather than querying the language model for each slot individually.
@@ A.1 Domain Classification
For the dialogue turns between the user and system, Which service is the user asking for? Choose from the following: restaurant, attraction, hotel, taxi, train, bus, hospital, police. Classify the user’s turn based on user intents, dialogue context, and previous turns. Don’t be biased toward domain-specific terms. If you can’t find a service, return None.
@@ Appendix B Language Models Hyperparameters
Appendix C LLM Request-Based Metrics for Resource Efficiency
We chose the number of prompt requests as a primary metric because it offers a consistent and practical means to evaluate resource efficiency across LLMs. Both time and cost are tightly linked to the frequency of API calls, particularly when using proprietary models where each request has an associated cost and latency. By focusing on the number of requests, we can bypass the complexities of traditional time complexity analysis, which is challenging to apply here given the black-box nature of these APIs. The lack of insight into system-level performance and hardware makes it impractical to assess time complexity in conventional terms, so tracking API requests provides a more straightforward and reliable measure of efficiency.
While prompt length is also a factor, it introduces additional variability that complicates analysis. Prompt length can depend on task details, slot values, descriptions, and dialogue structure, as well as API-specific requirements—some LLMs, for instance, necessitate resending the entire conversation history with each new prompt, while others do not. These differences further underscore the value of using the number of prompt requests as a central metric, allowing us to account for efficiency in a way that remains consistent regardless of such model-specific nuances.